### v2

# PlaceFM: A Training-free Geospatial Foundation Model of Places using Large-Scale Point of Interest Data

Abstract.
With the rapid growth and continual updates of geospatial data from diverse sources, geospatial foundation model pre-training for urban representation learning has emerged as a key research direction for advancing data-driven urban planning. Spatial structure is fundamental to effective geospatial intelligence systems; however, existing foundation models often lack the flexibility to reason about places, context-rich regions spanning multiple spatial granularities that may consist of many spatially and semantically related points of interest. To address this gap, we propose PlaceFM, a geospatial foundation model that captures place representations through a training-free, clustering-based approach. PlaceFM summarizes the entire point of interest graph constructed from U.S. Foursquare data, producing general-purpose region embeddings while automatically identifying places of interest. These embeddings can be directly integrated into geolocation data pipelines to support a variety of urban downstream tasks. Without the need for costly pre-training, PlaceFM provides a scalable and efficient solution for multi-granular geospatial analysis. Extensive experiments on two real-world prediction tasks, ZIP code–level population density and housing prices, demonstrate that PlaceFM not only outperforms most state-of-the-art graph-based geospatial foundation models but also achieves up to a 100× speedup in generating region-level representations on large-scale POI graphs. The implementation is available at https://github.com/mohammadhashemii/PlaceFM.
Geospatial Foundation Model, Graph-based Urban Representation Learning, General Purpose Encoders
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; August 31st-September 4th, 2026; Boston, MAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Geographic information systemsccs: Information systems Location based servicesccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representations
1. Introduction
The analysis of geospatial large datasets unlocks a wide range of applications that support diverse social functions across today’s world. In particular, recent research has harnessed the power of geospatial data to predict disease outbreaks (Kohn et al., 2023), detect behavioral patterns and anomalies (Amiri et al., 2024; Zhang et al., 2024a, b), monitor and forecast traffic, and inform smarter urban planning (Mokbel et al., 2024, 2022). Numerous machine learning methods have been developed across diverse data modalities, such as satellite imagery, search queries, economic indicators, and influenza trends, to learn informative representations of geospatial data (Du et al., 2020; Rolf et al., 2021; Ommi and Hashemi, 2024). In addition to these task-specific models, a range of prior approaches have also aimed to develop general-purpose geographic encoders that can be applied across multiple downstream geospatial tasks (Klemmer et al., 2025; Vivanco Cepeda et al., 2023). Despite their successes, existing machine learning approaches have often been limited by their task-specific nature, requiring carefully curated datasets and model architectures that generalize poorly across domains (Bommasani et al., 2021). This limitation has motivated the emergence of foundation models, large-scale pre-trained models that learn transferable representations from massive and diverse datasets. Foundation models first rose to prominence in natural language processing with the advent of large language models, and later transformed computer vision and multimodal models (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021). Their ability to generalize across tasks, modalities, and domains has led to widespread adoption in fields ranging from healthcare and biology to law and finance (Yang et al., 2022; Thirunavukarasu et al., 2023). Inspired by these advances, researchers have begun to explore geospatial foundation models and general-purpose encoders that aim to overcome the limitations of task-specific geospatial encoders (Liang et al., 2025).
Recently, several geospatial foundation models have been developed that focus on geolocation representation learning, aiming to generate general-purpose embeddings for geographic entities, such as urban regions (Liang et al., 2025; Tempelmeier et al., 2021; Li et al., 2022; Wu et al., 2023). Urban regions, as spatially distributed neighborhoods with relatively homogeneous physical and socioeconomic characteristics (Yuan et al., 2012), serve as a natural analytical unit for tasks such as region function recognition (Zhang et al., 2017, 2018), and population estimation (Cheng et al., 2022). By representing urban regions through embeddings learned from large-scale geospatial data, such as Points of Interest (POI) data or satellite images, these foundation models enable scalable and transferable analysis across multiple urban studies, bridging the gap between fine-grained location embeddings and aggregated urban patterns. In this context, it is reasonable to expect that learning meaningful representations for urban regions, capable of capturing and disentangling the underlying factors encoded in raw geolocation data, can fundamentally enhance a wide range of urban downstream tasks analyses (Mai et al., 2020).
To derive meaningful representations of urban regions from raw POI data, a common approach is to construct a POI graph in which nodes correspond to individual POIs and edges are defined using a geographical proximity function. General-purpose encoders are then trained on this graph to generate informative embeddings that capture the characteristics of urban regions(Jin et al., 2023; Yan et al., 2017; Agarwal et al., 2024; Huang et al., 2023). For instance, PDFM (Agarwal et al., 2024) pretrains on diverse geospatial data sources such as maps, activity levels, search trends, weather, and air quality by constructing a heterogeneous graph where nodes represent counties and postal codes and edges capture spatial proximity. A graph neural network (GNN) is then applied to learn embeddings for these geographic units. Similarly, HGI (Huang et al., 2023) aggregates POI data to the regional level with a multi-head attention mechanism, applies graph convolution to model spatial dependencies, and hierarchically combines these regional embeddings into city-level representations. Despite their potential, current geospatial foundation models for urban representation learning suffer from three key limitations that hinder their effectiveness in real-world applications:
(a) The Missing Sense of Place notion: While geospatial foundation models have advanced urban representation learning and demonstrated effectiveness across a variety of downstream tasks, current methods still struggle to capture the notion of places, regions shaped by human meaning and behavior that consist of spatially and semantically related POIs (Hashemi and Zufle, 2025). In other words, most existing methods produce learned representations at a fixed geographic granularity, such as ZIP codes, cities, and assign a single embedding to represent an entire region. The black-box process of generating a single vector to represent an entire region overlooks the fact that within any given area, multiple semantically meaningful subregions may exist that are highly relevant to users interests but not necessarily aligned with administrative boundaries. For example, a “cat lovers” place could emerge organically around a cluster of a cat-themed café, a park frequented by stray cats, and nearby pet shops, yet such meaningful patterns are lost when restricted to rigid spatial units. Moreover, representations derived solely from individual POIs, while useful as fine-grained descriptors of the built environment, fail to capture the broader semantic context that gives a region its true meaning (Niu and Silva, 2021).
(b) High Pre-training Computational Cost: A key challenge of existing geospatial foundation models for urban representation learning lies in the computational cost of the pre-training stage, which requires learning from massive POI datasets. Most current methods pre-train their encoders on a limited number of POIs, often by selecting sample cities or regions, which simplifies training but does not reflect real-world scales (Huang et al., 2023). Scaling to a nationwide dataset, for example, over twenty million POIs in the United States from Foursquare [^1] https://docs.foursquare.com/data-products/, would impose substantial computational overhead, demanding significant training time and resources while still aiming to learn effective and generalizable representations.
(c) Lack of Granularity Flexibility: Current geospatial foundation models are typically constrained to a single level of spatial granularity during both pre-training and inference time, restricting their flexibility in downstream applications. For example, PMT (Wu et al., 2024) encodes trajectories as sequences of United States Census Block Groups (CBGs), with all subsequent tasks, such as next-location prediction, restricted to that level. Similarly, PDFM (Agarwal et al., 2024) produces embeddings only for ZIP codes and counties in the United States, preventing inference at finer scales such as neighborhoods or CBGs. In practice, however, the notion of a place can be understood at multiple levels of granularity depending on the task and user perspective(Hashemi and Zufle, 2025) and may not align with political and administrative boundaries. For example, POIs related to a “cat lovers” place may span across multiple ZIP-code and CBG boundaries.
To address these three critical challenges, we propose PlaceFM, a training-free geospatial foundation model that is simple yet effective. PlaceFM is capable of (1) Multi-granular Region Representation Learning: generating meaningful representations for geographic entities across multiple levels of granularity, and (2) Place Identification: identifying places composed of spatially and semantically related POIs.
We begin by formally defining the notion of a place and outlining the motivation for identifying places within a geographical region composed of POIs. Subsequently, we introduce FSQ-19M, a large-scale POI graph dataset comprising approximately 19 million POIs constructed from the Foursquare tabular dataset, spanning all 48 contiguous U.S. states. Building on this resource, we propose a semantic-spatial-aware clustering-based approach for generating region embeddings while simultaneously identifying meaningful places within each region. Notably, PlaceFM enhances traceability by establishing an explicit correspondence between identified places and their original POIs, thereby providing deeper insights into how these places encapsulate information from the underlying POI graph. In summary, our contributions are:
  • We introduce FSQ-19M, a large-scale POI graph dataset with over 19 million POIs across the 48 contiguous U.S. states, constructed from the publicly available Foursquare dataset.
  • For the first time, we formalize the notion of a place in geospatial POI graph data and propose the first framework that automatically identifies places within a regional graph. The framework partitions POI representations into clusters, where the resulting centroids serve as place node embeddings.
[…]
In Section 2, we review existing geospatial foundation models and general-purpose encoders for urban representation learning, highlighting their key challenges and limitations. Section 3 then formalizes the problem setting, introduces the definition of a place, and outlines the scope of our study. The proposed methodology, including data integration strategies, model design, and system architecture, is detailed in Section 4. Section 5 presents the experimental setup, evaluation metrics, and performance results. Finally, Section 6 concludes the paper and outlines future research directions.
2. Related Works
2.1. Geospatial Foundation Models
Pre-trained geospatial foundation models are large-scale neural networks trained on diverse geospatial datasets, enabling them to capture transferable representations with broad utility across regions and tasks (Liang et al., 2025). Unlike traditional machine learning models that often fail to generalize beyond localized settings, these models can be effectively fine-tuned with limited task-specific data, offering scalability, adaptability, and efficiency for a wide range of spatial analysis applications. To address the issues in traditional machine learning approaches for geospatial data analysis, recent efforts have introduced geospatial foundation models designed for geolocation representation learning, with the goal of producing general-purpose embeddings for geographic entities (Liang et al., 2025; Agarwal et al., 2024; Balsebre et al., 2024; Tempelmeier et al., 2021; Li et al., 2022; Yan et al., 2024; Xiao et al., 2024; Wu et al., 2023). These embeddings serve as a backbone for a wide range of geospatial downstream tasks, facilitating more scalable and effective model development. For instance, GeoVectors (Tempelmeier et al., 2021) and SpaBERT (Li et al., 2022) leverage OpenStreetMap data [^2] https://www.openstreetmap.org/ to learn location embeddings, while G2PTL (Wu et al., 2023) is pre-trained on large-scale logistics delivery data.
Large-scale POI datasets have gained attention for pretraining geospatial foundation models(Jin et al., 2023). Being inherently multimodal, they combine signals such as satellite imagery, textual reviews, categorical attributes, and mobility patterns, enabling representations that capture both spatial and semantic properties. POI data can also be modeled as graph-structured data, where nodes represent locations and edges capture spatial or functional relationships (Chen et al., 2022). One of the closest related works, PDFM (Agarwal et al., 2024), frames geospatial pretraining as a graph learning problem. It constructs a heterogeneous graph where nodes represent counties and postal codes enriched with diverse signals such as maps, activity levels, search trends, weather, and air quality, while edges encode spatial proximity. A GNN is then employed to learn embeddings that capture complex geographic relationships. Despite the effectiveness of its embeddings in downstream tasks, PDFM is limited to generating general-purpose representations only at the ZIP code and county levels, restricting its ability to capture places at finer granularities such as neighborhoods or CBGs.
Figure 1. The spatial distribution of POIs based on a uniform random sample of one million entries from the FSQ-19M dataset, covering the 48 contiguous U.S. states.
2.2. Graph-based Urban Representation Learning
Urban representation learning aims to encode geographical entities, such as POIs and neighborhoods, into meaningful vector representations that capture both the functional and structural characteristics of cities. Such representations are crucial for understanding urban dynamics and supporting downstream tasks, including land use classification, location recommendation, and urban planning (Zhang et al., 2024c; Liang et al., 2025). Early approaches focused on learning embeddings from POIs to summarize the functional properties of geographical regions. Inspired by Word2Vec(Church, 2017), in (Yao et al., 2017a), embeddings for POI categories are learned based on co-occurrence patterns, while Place2Vec (Yan et al., 2017), improved this by leveraging K-nearest-neighbors (KNN) to capture contextual similarities between POIs. Subsequent works extended these methods to generate region-level representations using POI category embeddings(Zhai et al., 2019; Niu and Silva, 2021). However, these approaches often ignored the uniqueness of individual POIs: for instance, two restaurants in different spatial contexts would receive identical category embeddings, overlooking subtle yet meaningful differences. Recent studies addressed this limitation by incorporating the spatial context of each POI using GNN and message passing-based algorithms, defining region representation as a supervised graph classification problem (Xu et al., 2022). Modern urban representation learning methods have increasingly utilized graph-based approaches to capture relationships between regions. HGI (Huang et al., 2023) exemplifies this trend by aggregating POIs to the regional level using multi-head attention to model their diverse influences, applying graph convolution to encode similarities between adjacent regions, and further aggregating these regional representations into a city-level embedding. Such graph-centric models allow for more flexible and context-aware representations compared to methods restricted to fixed geographic units or category-level embeddings; however, realizing this potential, for example, in HGI, whose encoder is based on an attention-based neural network, requires training on massive POI datasets, incurring substantial computational overhead.
3. Preliminaries & Definitions
In this section, we provide a detailed description of all the notations and definitions:
  • $\mathcal{P}$: Set of all Points of Interest (POIs), where each POI $p\in\mathcal{P}$ is represented as $(\text{lat}_{p},\text{lon}_{p},A_{p})$, with $A_{p}=[a_{p_{1}},a_{p_{2}},\dots,a_{p_{K}}]$ denoting the list of $K$ attributes of $p$.
  • $a_{p_{i}}$: The $i$-th attribute of a POI $p$. In our study, we only use the hierarchical category attribute, though other attributes could be incorporated. As an example, $\text{category}_{p}$ as the hierarchical category attribute can be: ”Dining and Drinking $\rightarrow$ Middle Eastern Restaurant $\rightarrow$ Persian Cafe”.
  • $G=(\mathcal{V},\mathcal{E})$: Graph of POIs, where $\mathcal{V}=\mathcal{P}$ is the set of nodes, $\mathcal{E}$ is the set of edges, and edges are defined based on a proximity function $d(p_{i},p_{j})$.
  • $G_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$: Urban region, represented as a subgraph of POIs, with $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$.
3.1. Place Definition
Developing a geospatial foundation model requires moving beyond point-based map representations toward a structured concept of places (Hashemi and Zufle, 2025). Here, we define a place as a semantically meaningful spatial unit that may align with or encompass multiple geographic entities, including POIs, postcodes, neighborhoods, or administrative regions.
Definition 3.1 (Place).
A place $P$ is defined as a non-empty set of geographic entities:
$P=\{e_{1},e_{2},\ldots,e_{n}\},\quad\text{where }e_{i}\in\mathcal{E}$
$\mathcal{E}=P\cup\mathcal{P}$ denotes the universe of geographic entities, including primitive elements $\mathcal{P}$ (i.e., POIs) and higher-level places $P$ that are both semantically and spatially similar. This recursive definition allows us to capture the hierarchical nature of places, where smaller places can be nested within larger ones. The similarity between elements can be computed using any appropriate function, such as Euclidean distance or cosine similarity. The location of each entity $e_{i}$ is described by their coordinates $(\text{lat}_{e_{i}},\text{lon}_{e_{i}})$, where
[…]
3.2. Problem Statement
Let $G=(\mathcal{P},\mathcal{E})$ denote a graph of POIs, where $\mathcal{P}$ is the set of POIs and $\mathcal{E}$ encodes their spatial relationships. We assume a predefined set of disjoint urban regions $\mathcal{U}=\{G_{r_{1}},G_{r_{2}},\dots,G_{r_{N}}\}$, where $N=|\mathcal{U}|$ denotes the number of regions and each region, i.e., a subgraph $G_{r_{i}}\subseteq G$ corresponds to an administrative boundary such as a neighborhood, ZIP code, county, or city. Our problem consists of two main objectives:
3.2.1. Region-level representation learning
The first objective is to model urban environments across multiple spatial scales by compressing large-scale POI-level data into compact and informative region-level embeddings. Formally, for each region $G_{r_{i}}$ we seek to learn a $d$-dimensional feature vector $z_{r_{i}}\in\mathbb{R}^{d}$ such that
$z_{r_{i}}=f(G_{r_{i}}),$
where $f:\mathcal{G}\to\mathbb{R}^{d}$ is a representation function mapping graphs of POIs into a low-dimensional latent space. The generated embeddings $\{z_{r_{i}}\}_{i=1}^{N}$ should be both generalizable and effective across a wide range of downstream urban analytics tasks.
3.2.2. Place Discovery within Regions.
The second objective is to automatically identify higher-level places within each urban region $G_{r_{i}}$. A place $P\subseteq G_{r_{i}}$ as in Definition 3.1 is a subgraph consisting of POIs ro higher-level places that are both semantically and spatially similar. For each place $P$, we aim to compute an embedding
$z_{P}=f(P),$
such that $z_{P}$ captures the collective semantic and spatial characteristics of its constituent POIs. This hierarchical formulation ensures traceability, since we can explicitly associate each POI with its corresponding place, and compare similarity across places both within and across regions. Such capability has practical applications in place discovery (e.g., identifying emerging functional areas) and place recommendation (e.g., detecting multiple “cat-lovers” neighborhoods in a city).
Overall, the goal is to learn a representation function $f$ that supports hierarchical and multi-scale modeling of geographic space. This enables flexible analysis at different levels of granularity, from individual POIs to regions, while preserving the semantic interpretability and spatial coherence of discovered places.
4. Methodology
In this section, we provide a detailed overview of the data preparation process, the construction of the POI-level graph, and the explanation of the foundation model proposed in this paper. The architecture overview of our proposed method, PlaceFM, is depicted in Figure 2.
Figure 2. Pipeline of the proposed geospatial foundation model, PlaceFM. First, it builds POI-level graphs for each state, followed by category feature encoding and feature propagation to obtain neighborhood-aware POI embeddings. Places are then identified via training-free clustering at a chosen granularity, and an aggregator function produces the final region-level embedding.
4.1. FSQ-19M POI Dataset Preparation
Our FSQ-19M dataset consists of over 19 million preprocessed and enriched POIs across the entire contiguous United States, covering all 48 states. The data is extracted from the global Foursquare POI catalog which was downloaded from the official Foursquare Places platform [^3] https://foursquare.com/products/places/. Foursquare POI data has been extensively used in both industry and academic research (Zhao et al., 2016), as it provides high-quality, frequently updated, and semantically rich representations of locations such as restaurants, stores, parks, and service providers.
Pre-processing pipeline: After extracting all POIs located within the United States, we performed several pre-processing and enhancement steps to construct FSQ-19M. For each POI $p$, we retained its latitude ($\text{lat}_{p}$), longitude ($\text{lon}_{p}$), state, locality (city), and ZIP code. These attributes provide the spatial anchor for each point.
Table 1. A data entry in the FSQ-19M dataset.
Attribute Value
longitude 40.858911
latitude -124.073245
state CA
locality Arcata
postcode 95521
date_created 2012-02-08
date_closed None
category [ Dining and Drinking -> Restaurant -> Asian Restaurant -> Chinese Restaurant ]
Each POI is associated with the date it was created and, if applicable, the date it was closed, enabling temporal analysis of the evolution of urban spaces. Also, each POI is assigned a hierarchical semantic category, defined up to six levels. This hierarchical structure allows for flexible aggregation at different levels of semantic granularity. We applied preprocessing steps to normalize textual attributes (e.g., consistent naming of states and ZIP codes) and to remove duplicates. Moreover, we performed a spatial join with U.S. Census ZIP code boundary data to validate and, if necessary, correct the location of each POI. The enriched dataset covers diverse urban functions, dominated by dining venues such as pizzerias, coffee shops, and fast food restaurants. Fuel stations, hair salons, churches, and fire stations are also well represented, reflecting the dataset’s commercial and civic relevance for various spatial and socioeconomic analyses.
After preprocessing, the dataset consists of 19,019,187 POIs distributed across all 48 contiguous states in the U.S., each with spatial, temporal, and semantic information. Figure 1 visualizes the spatial distribution of POIs using a uniform random sample of one million entries from FSQ-19M. Also, Table 1 illustrates a representative sample from FSQ-19M, showing both spatial and semantic attributes. Section 4.2 describes the methodology used to construct a graph structure from the FSQ-19M POI dataset.
Table 2. General statistics of the FSQ-19M POI graph dataset
# Total Nodes (POIs) # Total Edges # Graphs (States) # Total regions (ZIP codes) # Total sub-regions (Localities)
19,019,187 44,202,952 48 31,970 60,155
Figure 3. Graph structures of a sample of POIs within ZIP code 10031 in Manhattan, NY: (a) 4-NN graph (unweighted), (b) 8-NN graph (unweighted), and (c) region-adaptive Delaunay triangulation (weighted). Edge opacity reflects weight, with darker edges indicating stronger connections.
4.2. POI Graph Data Construction
After obtaining the clean FSQ-19M POI-level data for the U.S., we construct 48 state-level geospatial graphs, each denoted as $G_{s}=(\mathbf{X}_{s},\mathbf{A}_{s})$, where $\mathbf{X}_{s}\in\mathbb{R}^{N_{s}\times d}$ represents the $d$-dimensional features of the $N_{s}$ POIs in state $s$ and $\mathbf{A}_{s}\in\mathbb{R}^{N_{s}\times N_{s}}$ is a weighted adjacency matrix encoding the spatial proximity between POIs. Constructing a graph is crucial because POIs with the same semantic features can have distinct functional roles depending on their surroundings. For instance, consider two coffee shops: one located in a busy shopping mall and another situated in a hospital complex. Despite sharing the same semantics, the former is likely to be associated with leisure and retail activity, whereas the latter may primarily serve healthcare visitors and staff. Capturing such contextual uniqueness is important for generating informative POI and region embeddings.
By representing POIs as nodes in a graph, each place, defined as a POI together with its surrounding spatial context, can be effectively modeled through the message-passing mechanism in GCNs. This allows each POI embedding to be enriched by propagating information from its neighboring POIs, thereby capturing the contextual semantics that shape the uniqueness of a place. The advantage of adopting a graph representation is that it provides a flexible and compact structure for modeling POIs, remains robust to spatial transformations, and naturally encodes contextual dependencies through message passing (Huang et al., 2023). Various strategies exist to construct such POI graphs, depending on the choice of proximity functions or similarity metrics for edge formation. In this study, we adopt two widely used strategies for constructing the graph structure based on geographic proximity.
(1) Region-Adaptive Delaunay Triangulation Delaunay triangulation (DT) (Delaunay et al., 1934) is a geometric algorithm that connects a set of points in the plane such that no point lies inside the circumcircle of any triangle in the triangulation. Many previous studies have verified the fitness of DT graphs for modeling the interactions among spatial vector data: this method ensures that edges connect spatially close POIs while avoiding overly dense connections, thereby preserving the local neighborhood structure while maintaining computational efficiency (Huang et al., 2023; Xu et al., 2022; Yan et al., 2019; Huang et al., 2022).
For each pair of POIs $i$ and $j$ connected by the triangulation, we assign a weighted edge in the adjacency matrix $\mathbf{A}_{s}$ defined as
$A_{ij}=\log\left(\frac{1+L_{r}^{1.5}}{1+d(i,j)^{1.5}}\right)\cdot w_{r}(i,j),$
where $d(i,j)$ denotes the geographic distance between POIs $i$ and $j$, which can be computed either as Euclidean distance in projected coordinates or Haversine distance on the sphere.
Here, $L_{r}$ is a region-specific scaling factor, computed using the POIs within the subgraph $G_{r}\subseteq G_{s}$ corresponding to a local region of POIs. This local, adaptive scaling is important because the spatial density of POIs can vary drastically even within the same state: for instance, POIs in a dense urban subarea such as Manhattan, NY, are much closer together than POIs in a less dense area of upstate New York, such as the Adirondack region. By computing $L_{r}$ for each region $G_{r}$, the weighting function remains balanced, ensuring that the edge weights are comparable across areas with highly variable local densities. Following (Huang et al., 2023), the factor $w_{r}(i,j)$ encodes regional consistency: it is set to $1.0$ if POIs $i$ and $j$ belong to the same region, and to $0.4$ if they connect across different regions, thereby down-weighting cross-region connections. The computed edge weights are then scaled uniformly within each region to the range $[0,1]$, resulting in $\mathbf{A}_{s}$ being a normalized weighted adjacency matrix.
(2) k-Nearest Neighbors (KNN). For each POI $i$, we connect its $k$ closest neighbors based on geographic distance (Euclidean or Haversine) in the adjacency matrix $\mathbf{A}_{s}$. Edges are unweighted (0 or 1), indicating the presence of a connection. Despite its simplicity, this approach balances neighbor consistency and spatial proximity: each node connects to its nearest neighbors, capturing local interactions while preventing overly dense graphs in high-density areas. Its efficiency and binary encoding make KNN graphs a practical choice for large-scale POI datasets and downstream GNN operations.
Figure 3 illustrates an example of different POI connectivity patterns in the 10031 ZIP code of Manhattan, NY. After constructing the graph structures for FSQ-19M, the statistics for all 48 graphs across the contiguous United States are summarized in Table 2.
4.3. Training-free Place Foundation Model
PlaceFM is a training-free foundation model of places, designed to generate meaningful urban region embeddings from graph-based POI data and to identify places within a given geographical region. The overall architecture of PlaceFM is shown in Figure 2. In this section, we provide a detailed description of each component of the framework:
4.3.1. Feature Encoder
Our proposed PlaceFM framework enables encoding POI-level raw attributes to extract rich and meaningful features, facilitating effective representation learning for a variety of urban-related tasks. In this study, we leverage the semantic category of each POI as the primary attribute, which is inherently hierarchical, capturing fine-grained distinctions between different types of POIs. To model these hierarchical category attributes, we utilize SD-CEM (Jia et al., 2024), a semantically disentangled POI category embedding model. It generates hierarchy-enhanced category representations by pre-training on large-scale mobility sequences, learning disentangled embeddings that capture semantic relationships among POI categories.
Formally, let $\mathcal{C}$ denote the set of POI categories with $L$ hierarchical levels. For a POI $p$, its category labels across levels are denoted by $\{c_{1},c_{2},\dots,c_{L}\}$, where $c_{1}$ is the most general level and $c_{L}$ the most specific. Formally. given a POI category $c_{i}$, it is mapped to a vector representation $\mathbf{v}_{c_{i}}\in\mathbb{R}^{d}$, such that
$\mathbf{v}_{p}=\text{SD-CEM}(\{c_{1},\dots,c_{L}\})\in\mathbb{R}^{d},$
where $\mathbf{v}_{p}$ captures both the hierarchical and semantic information of the POI.
4.3.2. POI Encoder
Upon encoding the features of each POI into a latent $d$-dimensional representation $\mathbf{v}_{p}$, we further need to enrich the representation of the POI itself. While the previous step captures the semantics of POI categories, it overlooks the uniqueness of each individual POI. In practice, POIs with the same semantic label can exhibit distinct characteristics depending on their surrounding environment. Intuitively, the uniqueness of a POI is shaped by its spatial context.
To capture the contextual semantics of places, it is therefore essential to move beyond isolated POI attributes and incorporate region-level structural information. For instance, a Starbucks located near university buildings and departments carries very different contextual meaning compared to one situated along a remote highway rest stop. To this end, PlaceFM employs a lightweight, non-parametric graph propagation mechanism that enriches each POI’s representation by incorporating information from its spatial neighbors.
Formally, let $G=(\mathcal{P},\mathcal{E})$ denote the POI graph, where $\mathcal{P}$ is the set of POIs and $\mathcal{E}$ the spatial edges connecting nearby POIs. Each node $p\in\mathcal{P}$ is initially associated with a feature vector $\mathbf{v}_{p}\in\mathbb{R}^{d}$. We follow the propagation method of Simplified Graph Convolution (SGC) (Wu et al., 2019), where the propagated representations after $k$ steps are computed as:
(2) $\mathbf{H}^{(k)}=\hat{\mathbf{A}}^{k}\mathbf{V},\quad\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times d},$
where $\hat{\mathbf{A}}$ is the symmetrically normalized adjacency matrix of $G$, defined as:
(3) $\hat{\mathbf{A}}=\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}},$
with $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ being the adjacency matrix with self-loops, and $\tilde{\mathbf{D}}$ the degree matrix of $\tilde{\mathbf{A}}$ such that $(\tilde{\mathbf{D}})_{ii}=\sum_{j}\tilde{\mathbf{A}}_{ij}$.
To aggregate information from multiple propagation depths, we adopt a weighted linear combination across steps:
(4) $\mathbf{H}=\sum_{k=0}^{K}\alpha_{k}\mathbf{H}^{(k)},$
where $\alpha_{k}\in\mathbb{R}$ is the propagation weight at step $k$. This formulation controls the extent to which semantic and spatial information is propagated, thereby allowing PlaceFM to act as a spatially and semantically controlled approach. In doing so, it ensures that each POI’s embedding integrates both local and multi-hop neighborhood semantics. Importantly, we allow $\alpha_{k}$ to vary across a wide range, including negative values. When $\alpha_{k}<0$, the model effectively introduces a “negative offset” that captures heterophilic relationships between POIs (Zhu et al., 2021), which is particularly useful in urban settings where co-located POIs may have contrasting functions (e.g., a bar next to a church).
4.3.3. Place Representation Learning via Clustering
Once we obtain the POI embeddings, the next step is to generate region-level embeddings and to identify places within a region, as introduced as the two main objectives of PlaceFM in Section 3. Unlike existing geospatial foundation models for urban representation learning, which are typically restricted to a fixed spatial resolution, PlaceFM is inherently multigranular, enabling the generation of embeddings at arbitrary levels of granularity. This flexibility accommodates both geographic zones (e.g., ZIP codes, cities, census block groups) and non-geographic groupings (e.g., functional regions, semantic clusters).
From a data-centric AI perspective (Zha et al., 2025), one effective approach to summarizing a large-scale POI graph while preserving its semantic and spatial structure is to learn representative embeddings for aggregated POIs within each region. This process, known as graph reduction (Hashemi et al., 2024), reduces the size of the graph while retaining sufficient expressivity for downstream tasks. Inspired by GECC (Gong et al., 2025), a clustering-based graph reduction method that guarantees comparable performance for GNNs trained on condensed graphs, PlaceFM introduces an urban representation learning component that leverages clustering over propagated POI embeddings to construct semantically meaningful and context-aware places. This transforms fine-grained POI embeddings into higher-level place representations.
Formally, let $\mathbf{H}\in\mathbb{R}^{n\times d}$ denote the matrix of POI embeddings for a region $r$, where $n$ is the number of POIs and $d$ the embedding dimension. For each predefined region $G_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$, we extract its embeddings:
(5) $\mathbf{H}_{r}=\{\mathbf{h}_{p}\mid p\in\mathcal{V}_{r}\},\quad\mathbf{H}_{r}\in\mathbb{R}^{n_{r}\times d},$
where $n_{r}=|\mathcal{V}_{r}|$ is the number of POIs in region $r$.
To extract places (i.e., sub-regions of semantically and spatially similar POIs), we apply bisecting $k$-means clustering (Rohilla et al., 2019). Unlike standard $k$-means, bisecting $k$-means iteratively splits the largest cluster into two sub-clusters until the desired number of clusters $k_{r}$ is reached, naturally producing a hierarchical structure that aligns with the hierarchical semantics of POI embeddings. The optimization objective is given by:
(6) $\min_{\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}}\sum_{j=1}^{k_{r}}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\|\mathbf{h}_{p}-\boldsymbol{\mu}_{j}\|_{2}^{2},$
where $\boldsymbol{\mu}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\mathbf{h}_{p}$ is the centroid of cluster $\mathcal{C}_{j}$.
Following GECC, we introduce a reduction ratio hyperparameter $r\in[0,1]$ to control the number of clusters per region: $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where smaller values of $r$ yield fewer but coarser places, and larger values of $r$ produce finer-grained places. This formulation allows flexible trade-offs between representational detail and computational efficiency. In cases where the distribution of POIs is highly non-uniform, density-based clustering methods such as DBSCAN (Ester et al., 1996) can also be employed, which do not require specifying $k_{r}$ in advance.
The resulting set of clusters $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k_{r}}\}$ serves as the places representing region $r$. Each cluster is summarized by its centroid embedding:
(7) $\mathbf{z}_{j}=\frac{1}{|\mathcal{C}_{j}|}\sum_{\mathbf{h}_{p}\in\mathcal{C}_{j}}\mathbf{h}_{p},\quad j=1,\dots,k_{r},$
yielding a condensed representation $\mathbf{Z}_{r}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k_{r}}\}\in\mathbb{R}^{k_{r}\times d}$. It has been theoretically proven (Gong et al., 2025) that such condensed representations are as expressive as the original POI embeddings, allowing downstream models trained on $\mathbf{Z}_{r}$ to achieve comparable performance to those trained on $\mathbf{H}_{r}$, while the size of $\mathbf{Z}_{r}$ is significantly smaller than $\mathbf{H}_{r}$, resulting in much higher efficiency during model training (Hashemi et al., 2024).
4.3.4. Region Aggregator
Once the place embeddings $\mathbf{Z}_{r}$ are obtained, in order to generate a single meaningful representation for the entire region $r$, we perform a simple and efficient aggregation function agg, which computes the weighted average of all place embeddings, weighted by the number of POIs in each place. Formally, the aggregated regional embedding is given by: $\mathbf{L}_{r}=\big[\,\mathbf{l}_{1}\;|\;\mathbf{l}_{2}\;|\;\dots\;|\;\mathbf{l}_{N_{r}}\,\big]$, where each $\mathbf{l}_{j}$ is a place embedding computed as the weighted average of POIs in place $\mathbf{l}_{j}=\frac{\sum_{i=1}^{N_{j}}n_{i}\cdot\mathbf{z}_{i}}{\sum_{i=1}^{N_{j}}n_{i}}$, where $\mathbf{z}_{i}$ is the embedding of place $i$, $n_{i}$ is the number of POIs contained in place $i$, and $N_{r}$ is the total number of places in region $r$. This weighted aggregation works because it ensures that places with more POIs, reflecting denser activity or higher functional importance, contribute proportionally more to the regional representation. In other words, the embedding $\mathbf{l}_{r}$ captures both the semantic characteristics of places and the relative concentration of POIs, leading to a more faithful and efficient summary of the entire region. If needed, we can also construct an adjacency matrix to represent the connectivity between region embeddings, where two regions are considered connected if their polygon boundaries touch each other.
5. Experiments
Table 3. ZIP-code level prediction performance for population density $\left(\text{people}/\text{km}^{2}\right)$. The best results are highlighted in bold, and values in parentheses indicate the standard deviation.
State (# POIs) Averaging Place2Vec (Yan et al., 2017) HGI (Huang et al., 2023) PDFM (Agarwal et al., 2024) PlaceFM (ours)
RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$
WY ($\sim 40K$) 65.67 22.44 -5.71 62.12 21.14 -1.22 69.52 19.66 -9.51 90.39 30.22 -13.56 61.14 18.14 1.27
($\pm$43.86) ($\pm$9.38) ($\pm$9.63) ($\pm$32.22) ($\pm$18.19) ($\pm$18.01) ($\pm$47.14) ($\pm$11.38) ($\pm$21.06) ($\pm$44.05) ($\pm$8.58) ($\pm$25.50) ($\pm$38.00) ($\pm$7.55) ($\pm$22.87)
VT ($\sim 52K$) 265.38 95.08 -13.83 240.76 78.08 -6.72 170.94 54.76 -1.33 308.18 107.07 -1.47 165.23 56.82 -0.89
($\pm$146.43) ($\pm$38.75) ($\pm$36.18) ($\pm$87.11) ($\pm$22.13) ($\pm$21.12) ($\pm$112.29) ($\pm$30.33) ($\pm$3.81) ($\pm$237.26) ($\pm$32.21) ($\pm$2.78) ($\pm$88.75) ($\pm$28.19) ($\pm$4.65)
AL ($\sim 255K$) 283.72 144.86 0.02 290.11 144.76 0.02 280.57 113.87 0.01 268.16 115.43 0.25 266.24 140.44 0.16
($\pm$27.90) ($\pm$12.13) ($\pm$0.48) ($\pm$22.32) ($\pm$16.12) ($\pm$0.29) ($\pm$46.08) ($\pm$10.77) ($\pm$0.62) ($\pm$68.00) ($\pm$11.01) ($\pm$0.17) ($\pm$18.83) ($\pm$9.31) ($\pm$0.35)
GA ($\sim 652K$) 532.00 197.31 0.36 511.21 187.14 0.26 458.21 167.16 0.53 536.34 186.33 0.38 410.16 155.02 0.89
($\pm$139.75) ($\pm$13.60) ($\pm$0.22) ($\pm$201.25) ($\pm$25.10) ($\pm$0.14) ($\pm$171.65) ($\pm$25.52) ($\pm$0.26) ($\pm$155.65) ($\pm$25.96) ($\pm$0.17) ($\pm$112.21) ($\pm$26.16) ($\pm$0.31)
NY ($\sim 1.24M$) 4872.13 1841.26 0.49 4644.87 1712.16 0.21 4305.66 1468.68 0.61 4483.95 1507.88 0.44 4410.86 1497.12 0.52
($\pm$722.78) ($\pm$246.76) ($\pm$0.10) ($\pm$589.28) ($\pm$126.06) ($\pm$0.12) ($\pm$843.69) ($\pm$270.21) ($\pm$0.09) ($\pm$700.85) ($\pm$300.87) ($\pm$0.11) ($\pm$642.54) ($\pm$318.22) ($\pm$0.12)
FL ($\sim 1.35M$) 1035.25 526.33 0.244 1115.62 615.21 0.16 772.30 351.21 0.50 806.30 373.98 0.56 702.56 310.72 0.35
($\pm$297.92) ($\pm$54.55) ($\pm$0.25) ($\pm$410.65) ($\pm$76.11) ($\pm$0.31) ($\pm$310.42) ($\pm$82.12) ($\pm$0.21) ($\pm$257.88) ($\pm$37.63) ($\pm$0.09) ($\pm$101.75) ($\pm$23.21) ($\pm$0.10)
CA ($\sim 2.2M$) 1665.47 963.78 0.40 1640.90 950.16 0.44 1540.18 819.69 0.50 1269.25 665.70 0.66 1053.43 588.14 0.69
($\pm$170.76) ($\pm$351.29) ($\pm$0.10) ($\pm$321.71) ($\pm$244.11) ($\pm$0.11) ($\pm$235.74) ($\pm$58.58) ($\pm$0.07) ($\pm$178.21) ($\pm$40.35) ($\pm$0.05) ($\pm$152.01) ($\pm$77.54) ($\pm$0.06)
To evaluate the effectiveness of PlaceFM, we compare it against the state-of-the-art baselines in urban representation learning. We begin by describing the experimental setup and implementation details of each method. We then report the performance of two urban downstream tasks: (i.) Population Density Prediction and (ii.) Housing Price Prediction, as measures of representation quality, along with a comparison of computational efficiency. Furthermore, we demonstrate the effectiveness of the identified places both quantitatively and qualitatively through case studies on selected regions using limited ground-truth data. Finally, we conduct an ablation study to empirically assess the contribution of each module in our framework.
5.1. Experimental Setup
5.1.1. Dataset and Baselines
The FSQ-19M POI graph dataset is constructed as a collection of 48 homogeneous graphs, each corresponding to one of the contiguous U.S. states. Splitting the dataset at the state level allows us to evaluate PlaceFM across graphs with diverse densities and structural statistics. For instance, Wyoming (WY) forms the smallest graph, containing 40,165 POIs, while California (CA) constitutes the largest, with 2,204,300 POIs. This setup provides a natural testbed for assessing the scalability and robustness of PlaceFM in comparison with existing methods under varying graph sizes and complexities.
We benchmark PlaceFM against several representative baselines, ranging from simple heuristics to state-of-the-art geospatial foundation models, all of which generate region-level embeddings within each state. For a fair comparison, the dimensionality of the generated region embeddings is fixed at $d=30$ across all models.
1. Averaging: A simple heuristic method in which the embedding of each region is computed by averaging the embeddings of its POI categories.
2. Place2Vec (Yan et al., 2017): A method that learns POI category embeddings and then derives region embeddings by averaging over the POIs contained in a region. Unlike simple averaging, Place2Vec incorporates spatial co-occurrence statistics during POI embedding learning, yet the final regional representation is still obtained through aggregation by averaging.
3. HGI (Huang et al., 2023): An unsupervised model that learns region embeddings by jointly modeling categorical semantics of POIs, POI-level and region-level adjacency, and the interaction between POIs and regions. HGI further employs an attention mechanism to weigh the relative importance of individual POIs during region-level aggregation, enabling more context-sensitive representations.
4. PDFM (Agarwal et al., 2024): A pre-trained foundation model that integrates diverse geospatial signals, including POI data, maps, activity levels, search trends, weather, and air quality, into a heterogeneous graph. Geographic regions such as counties and postal codes are represented as nodes, with edges reflecting spatial proximity. A GNN is then applied to capture spatial dependencies and generate embeddings for these regions.
5.1.2. Implementation Details
To ensure a fair reproduction and comparison with baseline methods, we reimplemented all baselines and tuned their hyperparameters, guided both by the best configurations reported in the original papers and by additional fine-tuning under our experimental setting. The complete implementation, including code for PlaceFM and all baseline models, is publicly available in the PlaceFM repository [^4] https://github.com/mohammadhashemii/PlaceFM. To maintain consistency, the number of evaluations for downstream tasks was restricted to 10. All baselines, with the exception of PDFM (Agarwal et al., 2024), were trained from scratch. For PDFM, due to the unavailability of detailed architectural specifications, we reused the released pretrained embeddings. Since PDFM integrates heterogeneous data sources (e.g., maps, activity levels, search trends, weather, and air quality), we only utilized the map-based embeddings and further applied dimensionality reduction to match the embedding size to $d=30$, ensuring a relatively fair comparison with other methods.
For PlaceFM, we determined that the optimal dimensionality of the feature encoder was $d=30$. The POI encoder was implemented using a single-layer SGC (Wu et al., 2019) with propagation limited to two steps (capturing up to second-order neighborhood information). The propagation coefficients were tuned within the range $\alpha_{0},\alpha_{1},\alpha_{2}\in[0.0,1.0]$ with increments of $0.25$. For the clustering module, the maximum number of iterations for $k$-means was set to 300, with a convergence threshold of $1\times 10^{-8}$. Each experiment was repeated 10 times, and we report the averaged results to ensure statistical reliability.
To efficiently execute the clustering procedure, we relied on Intel(R) Xeon(R) Platinum 8260 CPUs @ 2.40GHz with NumPy (Harris et al., 2020) for numerical computation. Baseline model training was conducted on a high-performance computing cluster equipped with a heterogeneous mix of GPUs: Tesla A100 (40GB) and V100 (32GB) for large-scale graphs, and K80 (12GB) GPUs for smaller graphs. This heterogeneous setup enabled efficient handling of datasets of varying sizes while ensuring consistent evaluation across all methods.
5.2. Downstream tasks performance analysis
5.2.1. Population Density Prediction
Understanding the spatial distribution of human populations is fundamental to a wide range of operational tasks, policy design, and scientific research, including disaster response, infrastructure development, and urban planning (Huang et al., 2023). Conventional census-based approaches to collecting population data, while reliable, are both labor-intensive and limited in spatiotemporal resolution. Consequently, recent studies have increasingly turned to alternative geospatial data sources, such as remote sensing imagery and POIs, to enable fine-grained population estimation (Yao et al., 2017b; Shang et al., 2021).
In this experiment, we leverage the region representations generated by PlaceFM to estimate population density at the ZIP-code level, which serves as a common unit of comparison with baseline methods. Although ZIP codes are chosen here for consistency, our framework is flexible and can be applied to regions of arbitrary granularity. Following baselines, to perform prediction, we train a Random Forest regressor (RF) with 100 decision trees, using 80% of the ZIP-code embeddings for training and the remaining 20% for testing. The ZIP-code level population density data used in our experiments was obtained from publicly available U.S. Census Bureau datasets.
To systematically assess scalability, we evaluate across states of varying sizes. For small-scale settings, we consider Wyoming (WY; 40,165 POIs across 166 ZIP codes) and Vermont (VT; 52,151 POIs across 250 ZIP codes). For medium-scale evaluation, we include Alabama (AL; 255,132 POIs across 626 ZIP codes) and Georgia (GA; 652,234 POIs across 719 ZIP codes). Finally, for large-scale experiments, we assess three diverse and dense states: New York (NY; 1,241,203 POIs across 1,738 ZIP codes), Florida (FL; 1,358,700 POIs across 973 ZIP codes), and California (CA; 2,204,300 POIs across 1,730 ZIP codes). This state selection captures a broad spectrum of graph sizes and densities, enabling a robust demonstration of PlaceFM’s scalability. The complete state-level results are available in the project repository.
Each experiment is repeated 10 times with randomized train-test splits, and average performance is reported. For evaluation, we adopt standard regression metrics: root mean squared error (RMSE), mean absolute error (MAE), and the coefficient of determination $(R^{2})$. Results are summarized in Table 3. For fairness, all baseline models are carefully reimplemented and tuned using their optimal hyperparameter configurations; these settings are reported in Table 4 for reproducibility.
Table 4. Optimal hyperparameter settings for each dataset. For each $\alpha_{i}$, the value shown to the left of the // corresponds to the Population Density prediction task, while the value on the right corresponds to the Housing Price prediction task.
POI Graph Reduction Ratio $r$ Clustering Method $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$
Wyoming (WY) 0.1 Kmeans 0.50.5/0.5 $0.5/0.0$ $0.0/0,0$
Vermont (VT) 0.1 Kmeans $0.0/0.0$ $0.5/0.25$ $0.0/0.0$
Alabama (AL) 0.1 Kmeans $0.0/0.0$ $0.0/0.25$ $0.5/0.0$
Georgia (GA) 0.05 Kmeans $0.5/0.5$ $1.0/1.0$ $0.0/0.0$
New York (NY) 0.02 Kmeans $0.25/0.75$ $0.0/0.0$ $0.0/0.0$
Florida (FL) 0.05 Kmeans $0.0/0.0$ $0.5/0.5$ $1.0/0.0$
California (CA) 0.05 Kmeans $0.0/1.0$ $0.5/0.5$ $1.0/1.0$
Empirical results demonstrate that PlaceFM consistently outperforms baseline approaches across nearly all datasets, underscoring the effectiveness of our training-free region representation learning paradigm. The only exception arises in New York, where PlaceFM ranks second by a marginal gap. We attribute this to the extreme density of the NY graph: when constructing the graph using the region-adaptive Delaunay Triangulation (DT) method, the resulting cluster representations tend to become homogenized, reducing discriminative capacity after POI encoding.
5.2.2. Housing Price Prediction
[…]
State (# POIs) Averaging Place2Vec (Yan et al., 2017) HGI (Huang et al., 2023) PDFM (Agarwal et al., 2024) PlaceFM (ours)
RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $R^{2}$ $\uparrow$
WY ($\sim 40K$) 370.45 217.81 -3.26 392.82 254.92 -4.02 361.39 220.70 -1.60 388.68 223.10 -3.17 350.21 202.88 2.87
($\pm$111.08) ($\pm$551.76) ($\pm$5.23) ($\pm$122.14) ($\pm$612.78) ($\pm$2.01) ($\pm$152.06) ($\pm$689.04) ($\pm$1.90) ($\pm$145.45) ($\pm$565.10) ($\pm$4.52) ($\pm$122.10) ($\pm$341.21) ($\pm$2.33)
VT ($\sim 52K$) 138.87 96.43 -0.20 144.02 111.38 -0.28 139.29 100.48 -0.23 210.18 157.12 -2.02 135.66 93.83 -0.16
($\pm$239.49) ($\pm$138.08) ($\pm$0.15) ($\pm$128.01) ($\pm$25.15) ($\pm$0.21) ($\pm$20.05) ($\pm$9.38) ($\pm$0.17) ($\pm$189.77) ($\pm$48.11) ($\pm$0.98) ($\pm$22.16) ($\pm$13.16) ($\pm$0.21)
AL ($\sim 255K$) 96.63 73.74 -0.17 96.57 72.16 -0.16 97.55 73.23 -0.20 96.14 72.38 -0.18 89.85 66.93 -0.02
($\pm$10.74) ($\pm$73.74) ($\pm$0.08) ($\pm$15.14) ($\pm$9.14) ($\pm$0.19) ($\pm$11.00) ($\pm$5.16) ($\pm$0.11) ($\pm$8.51) ($\pm$3.38) ($\pm$0.17) ($\pm$10.63) ($\pm$3.76) ($\pm$0.02)
GA ($\sim 652K$) 186.53 125.82 -0.24 192.01 131.20 -0.28 186.96 122.59 -0.25 197.52 118.50 -0.52 171.04 115.33 -0.02
($\pm$67.55) ($\pm$9.74) ($\pm$0.21) ($\pm$78.11) ($\pm$15.78) ($\pm$0.16) ($\pm$70.09) ($\pm$10.66) ($\pm$0.24) ($\pm$90.25) ($\pm$17.15) ($\pm$1.10) ($\pm$70.66) ($\pm$9.31) ($\pm$0.02)
NY ($\sim 1.24M$) 423.59 278.70 -0.10 421.65 270.21 -0.09 414.53 267.44 -0.06 410.11 251.08 -0.02 418.94 275.29 -0.07
($\pm$49.718) ($\pm$114.84) ($\pm$0.08) ($\pm$60.70) ($\pm$11.23) ($\pm$0.11) ($\pm$53.92) ($\pm$15.42) ($\pm$0.07) ($\pm$49.04) ($\pm$11.21) ($\pm$0.12) ($\pm$50.77) ($\pm$12.45) ($\pm$0.05)
FL ($\sim 1.35M$) 362.11 185.29 -0.20 382.28 201.07 -0.39 370.78 182.53 -0.19 375.30 189.67 -0.34 338.39 169.20 -0.01
($\pm$91.30) ($\pm$15.14) ($\pm$0.21) ($\pm$110.65) ($\pm$16.11) ($\pm$0.31) ($\pm$87.71) ($\pm$15.39) ($\pm$0.20) ($\pm$86.26) ($\pm$12.14) ($\pm$0.38) ($\pm$104.42) ($\pm$12.09) ($\pm$0.01)
CA ($\sim 2.2M$) 692.38 473.29 -0.08 688.21 467.00 -0.07 685.07 461.32 -0.06 701.09 471.23 -0.11 669.69 440.19 -0.01
($\pm$687.57) ($\pm$25.32) ($\pm$0.05) ($\pm$78.20) ($\pm$24.89) ($\pm$0.07) ($\pm$64.01) ($\pm$20.28) ($\pm$0.06) ($\pm$66.74) ($\pm$25.38) ($\pm$0.06) ($\pm$76.03) ($\pm$27.87) ($\pm$0.01)
Housing prices are a critical indicator of both social well-being and economic vitality, and their prediction has long been a central theme in urban studies, economics, and policy research. Accurate housing price estimation not only informs urban planning and economic forecasting but also guides individual decision-making in residential choices. Traditionally, two categories of factors are recognized as most influential: (i) structural attributes of properties, such as size, age, and physical condition, and (ii) locational amenities, which encompass access to services, infrastructure, and neighborhood characteristics. Within the latter category, POIs provide a particularly rich source of information, as they capture the functional and semantic attributes of urban spaces that directly influence housing demand and valuation (Huang et al., 2023).
In this experiment, we leverage the region-level embeddings generated by PlaceFM to predict housing prices at the ZIP-code level, thereby ensuring a consistent basis of comparison with baseline models. Following prior works, we adopt a Random Forest regressor (RF) with 100 decision trees, training on 80% of the ZIP-code embeddings while reserving the remaining 20% for testing. The housing price dataset is obtained from the Zillow Research data repository [^5] https://www.zillow.com/research/data/, specifically the Zillow Home Value Index (ZHVI) for August 2024, which provides a smoothed and standardized measure of typical home values across regions. To comprehensively assess scalability and robustness, we employ the same set of states as in the population density prediction task, encompassing graphs of varying sizes from small (e.g., Wyoming and Vermont) to large (e.g., New York, Florida, and California). Each experiment is repeated ten times with randomized train-test splits, and we report average performance to mitigate sampling variance. Again, for evaluation, we adopt standard regression metrics: RMSE, MAE, and $R^{2}$.
Experimental results, summarized in Table 5, demonstrate that PlaceFM consistently outperforms baseline methods across small- and medium-scale datasets. This highlights the effectiveness of our training-free approach in producing expressive region representations that generalize well across prediction tasks. A notable exception is observed in New York, where PlaceFM ranks third with only a marginal gap from the top-performing models. We believe that this gap arises for the same reasons as in the population density prediction task: the POI embeddings become overly homogenized during encoding, which diminishes their discriminative power.
Figure 4 illustrates the spatial distribution of absolute errors for Vermont and Georgia. The results show that predictions tend to be more accurate in city centers across both states, while larger discrepancies appear in peripheral and expansive regions. In Georgia, we also observe notable errors in highly dense urban areas, particularly in Atlanta. We attribute these patterns to several factors: First, larger regions tend to be more heterogeneous, making their internal variation harder to capture; Second, highly dense areas may exhibit complex housing market dynamics and sharp local variations that are difficult to model.
Figure 4. Spatial distribution of absolute housing price estimation errors. The top figure shows ZIP-code regions in Vermont (VT), and the bottom figure shows those in Georgia (GA).
5.3. Efficiency Comparison
Figure 5. Efficiency comparison of region embedding generation.
Figure 5 illustrates the efficiency and scalability advantages of PlaceFM, which primarily arise from its training-free embedding generation mechanism. For fairness, we exclude Averaging method and PDFM (Agarwal et al., 2024) from this comparison: Averaging incurs negligible computational cost as it simply computes the mean embedding for each region, whereas PDFM relies exclusively on pretrained embeddings from its original implementation. For all other methods, hyperparameters were selected to optimize performance on the downstream population density prediction task described in Section 5.2. For HGI (Huang et al., 2023), the number of training epochs was fixed at 100, fewer than those used in the original study.
As summarized in Figure 5, PlaceFM demonstrates consistent performance while efficiently managing computational resources even as graph size increases. For instance, in Wyoming, the embedding generation time for PlaceFM is 2.41 seconds for $r=0.1$ and 1.32 seconds for $r=0.05$, compared to 34.52 seconds for Place2Vec (Yan et al., 2017) and 60.61 seconds for HGI. Similarly, in Alabama, PlaceFM requires 14.61 seconds ($r=0.1$) and 7.81 seconds ($r=0.05$), while Place2Vec and HGI take 72.12 and 271.24 seconds, respectively. In Florida, the differences are even more pronounced, with PlaceFM completing in 94.65 seconds ($r=0.1$) and 48.29 seconds ($r=0.05$), versus 621.17 seconds for Place2Vec and 1293.84 seconds for HGI. These results indicate that PlaceFM achieves over an order of magnitude faster embedding generation while its runtime grows much more slowly with graph size.
5.4. Transferability Comparison
Table 6. Transferability: ZIP-code embeddings from PlaceFM with different regressor architectures. Units are RMSE, indicating the error of population density estimations (people/km2); lower is better. Best results are in bold.
Model Vermont (VT) Georgia (GA) Florida (FL)
RF MLP XGB RF MLP XGB RF MLP XGB
Averaging 265.38 321.45 315.64 532.00 612.11 576.42 1035.25 991.21 1141.67
Place2Vec (Yan et al., 2017) 240.76 246.20 242.12 511.21 498.23 488.25 1115.62 892.45 976.27
HGI (Huang et al., 2023) 170.94 182.55 175.98 458.21 423.05 413.01 772.30 762.43 730.67
PDFM (Agarwal et al., 2024) 308.18 292.11 254.21 536.34 500.07 478.68 806.30 714.09 726.89
PlaceFM (ours) 165.23 171.44 169.12 410.16 401.71 415.12 702.56 710.50 724.00
An essential criterion for assessing urban region representations is their ability to support diverse downstream architectures from a data-centric perspective. Unlike PlaceFM, which produces embeddings through a training-free mechanism, most existing approaches rely on backbone GNNs or other types of neural network architectures for embedding generation. This dependency can introduce inductive biases that limit their flexibility when applied to different downstream models.
Table 6 presents results demonstrating that embeddings derived from PlaceFM generalize robustly across three regression architectures: Random Forest (RF), Multi-Layer Perceptron (MLP, with two hidden layers of 32 and 16 neurons), and Extreme Gradient Boosting (XGB). With the exception of a single case in Georgia, where XGB trained on HGI embeddings achieves the best performance, PlaceFM consistently outperforms competing baselines. Reported values reflect the mean performance over ten runs of downstream inference, following the same setup as Section 5.2, with optimal hyperparameters selected for each model.
5.5. Multi-granular Place Identification
Figure 6. Voronoi spatial distribution of identified places in ZIP code 30329, Atlanta, GA. Places with the same color belong to the same cluster. For visualization clarity, only the top 10 identified places with the largest number of POIs are shown in each column. (a) Clustering using only category features. (b) and (c) Clustering using propagated category features.
Unlike existing geospatial foundation models, which primarily focus on urban region representation learning, PlaceFM pursues a second important objective: the automatic identification of places within each urban region. These places are defined as subgraphs consisting of POIs or higher-level clusters of POIs that are both semantically and spatially similar (see Definition 3.1). By learning to identify such places, PlaceFM enables the discovery of functionally coherent areas within an urban region, such as ZIP codes or cities. This is particularly useful for applications such as place recommendation, where users benefit from being directed toward clusters of POIs that match their preferences, and for urban analytics, where planners and policymakers can assess the functional value of different neighborhoods based on their composition of places (Islam et al., 2022; Zhao et al., 2016). For instance, identifying clusters of restaurants, cafés, and entertainment venues can highlight vibrant social districts, while clusters dominated by schools, libraries, and parks can indicate family-friendly neighborhoods.
Our clustering-based approach reduces the large-scale POI graph of each region into a predefined number of places, given by $k_{r}=\left\lfloor n_{r}\times r\right\rfloor$, where $n_{r}$ is the number of POIs in region $r$ and $r$ is the reduction ratio. The parameter $r$ determines the granularity of the identified places. Larger values of $r$ yield more fine-grained clusters, capturing smaller and more specialized neighborhoods (e.g., a cluster of restaurants and shops around a single intersection), whereas smaller values of $r$ lead to more general and aggregated places that encompass a wider variety of POI types (e.g., a commercial district spanning multiple blocks). In this way, $r$ serves as a natural control parameter for the level of granularity at which urban functionality is represented.
Figure 6 illustrates this process using Voronoi diagrams (Aurenhammer and Klein, 1996), which depict the nearest-area partitioning around each POI in ZIP code 30329, Atlanta, GA, after the place generation process. POIs sharing the same Voronoi color are clustered into the same place. The upper row provides a $3\times$ zoomed view of the lower row to highlight POIs in greater detail. For visualization purposes, in all three cases we display only the top 10 identified places with the largest number of POIs after clustering:
Semantic-only-based Places. In column (a), representations of places are obtained by clustering solely on semantic information, using only category-level POI features. This results in discontinuous and fragmented clusters, where each color corresponds to a semantic category of POIs. For instance, the dark blue cluster represents POIs belonging to category level 1 of Dining and Drinking, which typically includes restaurants, cafés, bars, and similar establishments. This approach overlooks the influence of neighboring POIs in the vicinity of each POI, thereby ignoring valuable spatial context when constructing a representation for a place.
Semantic- and Spatial-Based Places. Unlike in column (a), in columns (b) and (c), PlaceFM leverages both semantic and spatial feature aggregation, producing spatially contiguous and semantically coherent clusters that better reflect meaningful urban places. This property is particularly useful for the traceability of how place embeddings are generated: First, each embedding can be directly linked back to the underlying POIs that constitute the place, along with their spatial configuration and semantic similarity. Such traceability ensures that the learned representations are interpretable, allowing us to understand why two regions are considered functionally similar or different. Second, by aggregating POIs into coherent clusters, we obtain a structured view of the urban landscape, where the relationship between individual places and the overall region becomes explicit. This facilitates not only better transparency in the embedding generation process but also a richer understanding of the functional organization of an entire region.
In column (c), the places with red Voronoi color contain mostly POIs categorized under Dining and Drinking, along with some from Business and Professional Services. This indicates areas that function as mixed-use clusters, where restaurants, cafés, and bars are co-located with offices or service-oriented businesses, reflecting the multifunctional character of urban neighborhoods. In real-world settings, such patterns are commonly found in downtown districts or commercial corridors of large cities, where dining establishments are interwoven with professional offices, coworking spaces, and service providers, creating vibrant hubs of both social and economic activity. .
Effect of reduction ratio $r$. As the reduction ratio $r$ increases, more clusters emerge, capturing finer-grained structures, while smaller values of $r$ highlight broader, more general areas. For example, by doubling $r$, the number of identified places also roughly doubles, but the composition of POIs within each place becomes more unique and specialized. In column (b), compared to column (c), identified places appear more discontinuous and fragmented; however, each place highlights semantically and spatially coherent POIs in greater detail. For instance, in column (b), the green Voronoi areas contain POIs predominantly from category level 1 of Community and Government and Business and Professional Services, which may represent local government offices, community centers, libraries, as well as law firms or consulting offices. Such detailed, specialized clusters are less apparent in the more general, aggregated places shown in column (c), demonstrating how varying $r$ allows control over the granularity of identified urban places and their functional specificity.
6. Conclusion and Future Work
We introduced PlaceFM, a geospatial foundation model that generates general-purpose place embeddings by capturing spatial context and neighborhood structure. Our comprehensive experiments demonstrate its effectiveness in producing region-level embeddings at multiple geographic scales, which can be readily applied to a variety of urban downstream tasks. In this study, the only data modality used for POI features was category information. As future work, PlaceFM could be extended to leverage multi-modal data, including mobility information, to capture human movement patterns, enabling the learning of more informative and meaningful region-level and place-level representations.